CAT Chat Minutes

October 22, 1999

Information and Follow-up Items

1. A power outage took place on Saturday, October, 16. What happened?

A switchgear supplying power to the injector was shut down for service. As it was shut down, the main switchgear faulted for no apparent reason. The recovery was quick, and the cause is being investigated.

2. A while ago, it had been mentioned that dosimeters may not be needed on the experiment hall floor. Are there any updates?

Up to this point, there have been no recordable doses on the experiment hall floor. However, we will continue requiring badges on the floor as we build up better statistics for the various operating modes.

3. Is there a way to improve the water temperature control in the LOM showers? It seems to take quite a bit of time to get hot water.

We have uncovered some plumbing problems and will begin correcting them.

4. The LOM vending machines are always empty. Is there a way we can get the vendor to correct this?

PFS has been contacted and has been made aware (again) of the supply problem.

User Operations

This is the last week of the shutdown. Machine studies will begin Friday, October 29 at 0800 hours. Shortly thereafter, User Operations will begin on Thursday, November 4 at 0800 hours.

General Information

Gopal Shenoy is scheduled to go home today from the hospital. We are asking everyone not to call or visit him until he is stronger.

All maintenance activities are proceeding on schedule. A minor setback occurred when a water leak was discovered on a photon shutter in the 17-ID front end. The component was replaced, and no impact on the schedule is expected. The steam shutdown on Thursday for a valve replacement was completed without any problems.

The liquid nitrogen system installation is on schedule, and the installed systems are working well.

The web page showing beam parameters for a 24-hour period is being finalized and will be made available for this run.